### v10

# Movement Prediction-Adjusted Naive Forecast: Is the Naive Baseline Unbeatable in Financial Time Series Forecasting?

Abstract
In financial time series forecasting, the naive forecast is a notoriously difficult benchmark to surpass because of the stochastic nature of the data. Motivated by this challenge, this study introduces the movement prediction-adjusted naive forecast (MPANF), a forecast combination method that systematically refines the naive forecast by incorporating directional information. In particular, MPANF adjusts the naive forecast with an increment formed by three components: the in-sample mean absolute increment as the base magnitude, the movement prediction as the sign, and a coefficient derived from the in-sample movement prediction accuracy as the scaling factor. The experimental results on eight financial time series, using the RMSE, MAE, MAPE, and sMAPE, show that with a movement prediction accuracy of approximately 0.55, MPANF generally outperforms common benchmarks, including the naive forecast, naive forecast with drift, IMA(1,1), and linear regression. These findings indicate that MPANF has the potential to outperform the naive baseline when reliable movement predictions are available.
Keywords Financial time series $\cdot$ Point forecasting $\cdot$ Movement prediction $\cdot$ Naive baseline $\cdot$ Forecast combination
1 Introduction
Financial time series forecasting plays a crucial role in many decision-making processes. The persistent demand for accurate point forecasts in financial markets has drawn sustained attention from researchers and practitioners. However, although financial time series have long served as testbeds for forecasting methods, their stochastic nature continues to pose a challenge in achieving accurate predictions (Hamilton,, 2020). One prominent theory on this challenge is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970). The naive forecast, which predicts the next value of the target variable to be the same as its current value, is frequently regarded as the optimal benchmark for these tasks.
In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naive forecast. Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naive forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite numerous attempts to exploit weak structural signals in financial time series, existing forecasting methods, including statistical, machine learning, and deep learning approaches, have typically failed to outperform the naive baseline (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
From the forecaster’s perspective, the apparent unpredictability of a variable may stem from an incomplete understanding of the underlying processes, limited knowledge of contributing factors, or limitations in existing analytical methods (Laplace,, 2012). In this sense, randomness in data may not be entirely inherent but may be partly a reflection of current modeling constraints (Crutchfield and Feldman,, 2003). While useful information can be present in the input space, it is often masked by irrelevant signals, which makes it difficult for regression methods to fully capture the informative content (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005). Furthermore, identifying the predictors that are most relevant to the target variable remains a complex task (Granger,, 1969; Stock and Watson,, 2002). Nevertheless, these perspectives and empirical findings, while highlighting the practical difficulties of uncovering exploitable patterns, do not preclude the predictability that may exist in the data.
From an information-theoretic perspective, forecasting is theoretically possible when the mutual information between the predictors and the target variable is greater than zero, as this confirms that a nonzero dependency exists (Kraskov et al.,, 2004; Cover and Thomas,, 2012). However, practical forecasting feasibility depends on the magnitude, nature, and stability of this mutual information. In binary forecasting tasks, such as predicting whether a value will rise or fall, even modest dependencies can reduce uncertainty and enable models to outperform random guessing (Ircio et al.,, 2020). In contrast, for continuous point forecasting tasks, a similar magnitude of dependency may only marginally reduce the overall uncertainty of the numerical output (Beraha et al.,, 2019). This asymmetry partly accounts for the abundance of research on movement prediction in financial time series.
Indeed, previous studies have noted that movement prediction is generally more tractable than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013). Behavioral finance studies suggest that investor sentiment and psychological factors can generate predictable movement patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Barberis et al.,, 1998; Baker and Wurgler,, 2007). Consequently, predictors such as sentiment indicators are often used to improve movement prediction accuracy. For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 movement prediction accuracy. Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, and the movement prediction accuracy reached 0.70. Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short-term memory (LSTM) network and a transformer for stock movement prediction, and the accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the movement prediction accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
Since various studies have demonstrated the feasibility of meaningful movement prediction, which can provide directional information that the naive forecast cannot offer, it is natural to wonder whether such directional information can be exploited to improve the point forecast performance. Guided by the “less is more” principle, which suggests that reducing informational complexity can improve decision-making (Gigerenzer and Brighton,, 2009), a method that aims to incorporate directional information for point forecasting can rely solely on naive forecasts and movement predictions as inputs and can be considered a combination of the two types of forecasts. In doing so, the method avoids the potential complications introduced by additional predictors that may harm the final point forecast performance and can serve as the final step in a forecasting framework, as shown in Figure 1. Unfortunately, existing research has largely treated movement prediction and point forecasting as separate tasks, and most forecast combination methods focus on aggregating similar types of forecasts. To the best of our knowledge, few studies have examined how and why movement prediction can be integrated with naive forecasts to potentially improve point forecasting performance.
Figure 1: Forecasting framework incorporating naive forecast and movement prediction.
To investigate the potential of combining the movement prediction with the naive forecast for financial time series, this study proposes a forecast combination method, named the movement prediction-adjusted naive forecast (MPANF). In particular, MPANF can use the output of any classification model and adjust the naive forecast with an increment formed by three components: the in-sample mean absolute increment as the base magnitude, the movement prediction as the sign, and a coefficient derived from the in-sample movement prediction accuracy as the scaling factor. The effectiveness of MPANF is evaluated empirically on eight real-world financial time series from the U.S. market. In all the evaluations, the naive forecast, naive forecast with drift, integrated moving average of order (1,1) (IMA(1,1)), and linear regression serve as baselines, with the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE) and symmetric mean absolute percentage error (sMAPE) used as evaluation metrics.
This study makes four significant contributions to the literature:
  • First, it introduces a novel forecast combination method that systematically incorporates movement prediction into point forecasting, thereby linking two tasks that have typically been studied separately. While binary classification results have previously been used to guide regression tasks, using the target variable’s own movement prediction as the sole complementary information to enhance the naive forecast provides a fresh methodological perspective.
  • Second, it provides a theoretical foundation for why combining movement prediction with a naive forecast is superior to using the naive forecast alone. The analytical estimate of the optimal coefficient establishes a connection between movement prediction accuracy and improvement upon the naive forecast, underscoring the informational value of directional signals and distinguishing the proposed method from previous forecasting methods.
  • Third, it derives a retrospective condition, the joint requirement on the out-of-sample coefficient and mean absolute increment, for MPANF to outperform the naive baseline. This condition provides both a theoretical basis and a diagnostic tool, moving beyond the common practice of explaining forecasting performance merely as “overfitting” or “good generalization” by offering a concrete, testable criterion.
  • Fourth, it provides empirical evidence on the conditional predictability of financial time series. For time series that are often regarded as challenging to predict, MPANF can outperform the naive baseline when accurate movement predictions are available.
The remainder of this paper is organized as follows. Section 2 presents the details of MPANF and explains its theoretical framework and algorithmic implementation. Section 3 reports experimental results on its performance using real-world financial time series. Section 4 discusses the implications of the findings and the limitations of this study. Finally, Section 5 concludes this study and outlines directions for future research.
2 MPANF
In this section, we first outline the MPANF’s general formulation, followed by the analytical estimate of its optimal coefficient and the retrospective condition that governs its empirical performance. Next, we detail the algorithmic implementation of the MPANF for out-of-sample forecasting.
2.1 General Formulation
The value of a time series at time step $t$ can be expressed as the sum of its previous value at time step t1t-1 and an increment ${\epsilon}_{t}$} between two adjacent observations. This relationship can be expressed as:
${{y}_{t}}={{y}_{t-1}}+{{\epsilon}_{t}},$ (1)
where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step t1t-1. The increment $\epsilon_{t}$ at time step $t$ can be expressed as a product of two terms: a nonnegative term $|\epsilon_{t}|$ representing the magnitude of the increment and a sign term ${d}_{t}$, which indicates the direction of the increment between two adjacent points and takes a value of +1+1 for upward movement and 1-1 for downward movement. This allows ${y}_{t}$ to be reformulated as follows:
${{y}_{t}}={{y}_{t-1}}+{{d}_{t}}\cdot|\epsilon_{t}|;|\epsilon_{t}|\geq 0$ (2)
The nonnegative term $|\epsilon_{t}|$ can be further decomposed into two components: $\bar{\epsilon}$, representing the mean absolute increment of the time series, and $\alpha_{t}$, a nonnegative coefficient that controls the magnitude of the increment at time step $t$. Accordingly, ${y}_{t}$ can be expanded as follows:
${{y}_{t}}={{y}_{t-1}}+{{d}_{t}}\cdot{\alpha_{t}}\cdot\bar{\epsilon};{\alpha_{t}}\geq 0,\bar{\epsilon}>0,$ (3)
Eq. (3) implies that the point forecast of the target variable can be generated if ${d}_{t}$} can be predicted and $\alpha_{t}$ can be estimated. Consider a simplified forecasting strategy in which the increments for each future time step are set to have a fixed magnitude; in other words, we set $\alpha_{t}$ to be a constant value and derive the one-step-ahead point forecast at time step $t$, denoted by $\hat{y}_{t}$, as follows:
${{\hat{y}}_{t}}={{y}_{t-1}}+\hat{d}_{t}\cdot\alpha\cdot\bar{\epsilon};\bar{\epsilon}>0,$ (4)
where ${\hat{d}}_{t}$} is the movement prediction, $\alpha$ is a fixed scalar, and $\alpha\in\mathbb{R}$. According to Eq. (4), $\hat{y}_{t}$ can be regarded as an adjustment of the naive forecast with an increment formed by three components: $\bar{\epsilon}$ as the base magnitude, ${\hat{d}}_{t}$} as the sign, and $\alpha$ as the scaling factor. As movement prediction plays a central role in enabling this adjustment, the resulting forecast is referred to as the movement-prediction-adjusted naive forecast (MPANF).
2.2 Optimal Coefficient
Given a time series and its corresponding movement prediction, only the coefficient $\alpha$ in the MPANF requires optimization. Suppose that the mean squared error (MSE) is used as the cost function; then, the original optimization problem can be replaced by an equivalent problem, which can further be approximated as a convex surrogate. Solving this convex approximation yields a closed-form expression for $\alpha$, which we interpret as an analytical estimate of the optimal in-sample coefficient for the MPANF.
As the in-sample MSE of the naive forecast is a constant value, minimizing the in-sample MSE of the MPANF is equivalent to maximizing the in-sample MSE difference between the naive forecast and the MPANF:
$minimize\ MSE_{\text{in}}^{\text{MPANF}}\equiv\ maximize\left(MSE_{\text{in}}^{\text{Naive}}-MSE_{\text{in}}^{\text{MPANF}}\right),$ (5)
where $MSE_{\text{in}}^{\text{Naive}}$ and $MSE_{\text{in}}^{\text{MPANF}}$ denote the in-sample MSEs of the naive forecast and the MPANF, respectively. Suppose that the in-sample set has $N$ time steps and that the in-sample movement prediction yields $n_{\text{in}}^{\text{correct}}$ correct and $n_{\text{in}}^{\text{incorrect}}$ incorrect directions; then, the in-sample movement prediction accuracy, denoted by $ACC_{\text{in}}$, can be expressed as follows:
$ACC_{\text{in}}=\frac{n_{\text{in}}^{\text{correct}}}{n_{\text{in}}^{\text{incorrect}}+n_{\text{in}}^{\text{correct}}}$ (6)
Assuming that the in-sample movement prediction is unbiased and contains sufficient numbers of both correct and incorrect predictions, the in-sample MSE difference between the naive forecast and the MPANF, denoted by $\Delta MSE_{\text{in}}$, can be approximated as follows:
$\Delta MSE_{\text{in}}$ $=MSE_{\text{in}}^{\text{Naive}}-MSE_{\text{in}}^{\text{MPANF}}$ (7)
$\approx(4\cdot\alpha_{\text{in}}\cdot ACC_{\text{in}}-\alpha_{\text{in}}^{2}-2\cdot\alpha_{\text{in}})\cdot\bar{\epsilon}_{\text{in}}^{2},$
where $\alpha_{\text{in}}$ denotes the in-sample coefficient $\alpha$ and $\bar{\epsilon}_{\text{in}}$ denotes the in-sample mean absolute increment. A detailed algebraic derivation of the $\Delta MSE_{\text{in}}$ is provided in Appendix A.
According to Eq. (7), maximizing the approximation of $\Delta MSE_{\text{in}}$ is equivalent to maximizing the following concave function of $\alpha_{\text{in}}$:
$f(\alpha_{\text{in}})=4\cdot\alpha_{\text{in}}\cdot ACC_{\text{in}}-\alpha_{\text{in}}^{2}-2\cdot\alpha_{\text{in}}$ (8)
By differentiating this function with respect to $\alpha_{\text{in}}$ and setting the derivative to zero, we obtain the analytical estimate of the optimal in-sample coefficient, which is formalized in Proposition 1.
Proposition 1 (Coefficient Estimate).
The analytical estimate of the optimal in-sample coefficient, denoted by $\hat{\alpha}_{\text{in}}^{*}$, which is the optimal solution for a convex approximation of the MSE difference, is given by:
$\hat{\alpha}_{\text{in}}^{*}=2\cdot ACC_{\text{in}}-1$ (9)
Eq. (9) indicates that this estimate can be derived using only the in-sample movement prediction accuracy and is independent of the characteristics of the in-sample series. Substituting this analytical estimate from Eq. (9) back into the approximated MSE difference function based on Eq. (7) yields the theoretical in-sample MSE improvement, as formalized in Proposition 2.
Proposition 2 (MSE Improvement).
Using the analytical estimate of the optimal in-sample $\alpha$, the MPANF theoretically achieves a nonnegative MSE improvement over the naive forecast on the in-sample set:
$\max\Delta MSE_{\text{in}}\approx\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\geq 0$ (10)
As Proposition 2 demonstrates, when a movement prediction is better than random guessing, a measurable in-sample MSE improvement by the MPANF over the naive baseline is achievable. This improvement is proportional to the movement prediction accuracy: as the accuracy increases, the in-sample MSE improvement also increases. At a movement prediction accuracy of 1.0, the maximum achievable MSE improvement is approximately $\bar{\epsilon}_{\text{in}}^{2}$. Thus, under idealized conditions, the MPANF’s expected performance in terms of the MSE matches or surpasses that of the naive forecast on the in-sample set:
$MSE_{\text{in}}^{\text{MPANF}}\leq MSE_{\text{in}}^{\text{Naive}}$ (11)
2.3 Retrospective Condition
Suppose that $\hat{\alpha}_{\text{in}}^{*}$ and $\bar{\epsilon}_{\text{in}}$ are adopted by the MPANF for the out-of-sample forecast and that the out-of-sample set contains $M$ time steps and the out-of-sample movement prediction yield $m_{\text{out}}^{\text{correct}}$ correct and $m_{\text{out}}^{\text{incorrect}}$ incorrect predictions; then, the out-of-sample movement prediction accuracy, denoted by $ACC_{\text{out}}$, can be expressed as:
$ACC_{\text{out}}=\frac{m_{\text{out}}^{\text{correct}}}{m_{\text{out}}^{\text{incorrect}}+m_{\text{out}}^{\text{correct}}}$ (12)
The out-of-sample MSE difference between the naive forecast and the MPANF, denoted by $\Delta MSE_{\text{out}}$, can subsequently be approximated as follows:
$\Delta MSE_{\text{out}}$ $=MSE_{\text{out}}^{\text{Naive}}-MSE_{\text{out}}^{\text{MPANF}}$ (13)
$\approx 2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot\hat{\alpha}_{\text{out}}^{*}\cdot\bar{\epsilon}_{\text{out}}-\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2},$
where $\bar{\epsilon}_{\text{out}}$ denotes the out-of-sample mean absolute increment and $\hat{\alpha}_{\text{out}}^{*}$ denotes the analytical estimate of the optimal out-of-sample coefficient, which is derived according to Proposition 1:
$\hat{\alpha}_{\text{out}}^{*}=2\cdot ACC_{\text{out}}-1$ (14)
A detailed algebraic expansion of the $\Delta MSE_{\text{out}}$ is provided in Appendix B. For the MPANF to outperform the naive baseline in terms of the out-of-sample MSE, the approximated $\Delta MSE_{\text{out}}$ should be no less than zero, which leads to a retrospective condition for out-of-sample outperformance, as formalized in Proposition 3.
Proposition 3 (Retrospective Condition).
The retrospective condition for the MPANF to achieve a lower out-of-sample MSE than the naive baseline is given by:
$\hat{\alpha}_{\text{out}}^{*}\cdot\bar{\epsilon}_{\text{out}}\geq\frac{1}{2}\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}$ (15)
Since $\hat{\alpha}_{\text{out}}^{*}$ and $\bar{\epsilon}_{\text{out}}$ cannot be determined in advance, we can treat them as two random variables, each following a distribution defined over nonnegative values. Accordingly, Eq. (15) represents a probabilistic event whose likelihood depends on the realized values of these two variables. Smaller values of these two variables generally increase the likelihood of improvement. However, as shown in Eq. (13), the resulting improvements may be marginal. The MPANF can realize its theoretical maximum improvement in terms of the out-of-sample MSE only when $\hat{\alpha}_{\text{in}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{in}}$ are close to $\hat{\alpha}_{\text{out}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out}}$, respectively.
2.4 Algorithmic Implementation
The MPANF is a versatile forecast combination method that can use movement predictions from any external model. This modularity allows the MPANF to integrate advancements in time series movement prediction without altering its core framework. If the movement prediction is provided, the practical implementation of the MPANF in a one-step-ahead scenario is summarized in Algorithm 1.
Algorithm 1 Movement Prediction–Adjusted Naive Forecast
procedure MPANF($Y_{1:N},\ \hat{d}^{\text{in}}_{2:N},\ \{\hat{d}^{\text{out}}_{t+1}\}_{t\geq N}$)
  $P\leftarrow[\,]$
  $\bar{\epsilon}_{\text{in}}\leftarrow\dfrac{1}{N-1}\sum_{t=2}^{N}\lvert Y_{t}-Y_{t-1}\rvert$
  $d^{\text{in}}_{t}\leftarrow\mathrm{sign}(Y_{t}-Y_{t-1})\ \ \text{for }t=2,\dots,N$
  $ACC_{\text{in}}\leftarrow\dfrac{1}{N-1}\sum_{t=2}^{N}\mathbf{1}\!\big(\hat{d}^{\text{in}}_{t}=d^{\text{in}}_{t}\big)$
  $\hat{\alpha}_{\text{in}}^{*}\leftarrow 2\cdot ACC_{\text{in}}-1$
  $Y_{\mathrm{last}}\leftarrow Y_{N}$; $t\leftarrow N$
  while next $\hat{d}^{\text{out}}_{t+1}$ is available do
   $\Delta\leftarrow\hat{d}^{\text{out}}_{t+1}\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}$
   $\hat{Y}_{t+1}\leftarrow Y_{\mathrm{last}}+\Delta$
   append $\hat{Y}_{t+1}$ to $P$
   observe $Y_{t+1}$
   $Y_{\mathrm{last}}\leftarrow Y_{t+1}$; $t\leftarrow t+1$
  end while
  return $P$
end procedure
Algorithm 1 first computes the in-sample mean absolute increment, $\bar{\epsilon}_{\text{in}}$, and the in-sample movement prediction accuracy, $ACC_{\text{in}}$, from the provided in-sample data. These values are then used to calculate the analytically derived coefficient, $\hat{\alpha}_{\text{in}}^{*}$. In the forecasting loop, for each new time step, the algorithm generates a forecast, $\hat{Y}_{t+1}$, by adjusting the most recent observation, $Y_{\text{last}}$. This adjustment, $\Delta$, is the product of the movement prediction on $t+1$, the coefficient, and the in-sample mean absolute increment. Through this procedure, the algorithm accumulates the set of out-of-sample forecasts.
3 Experiments
This section details the experiments for evaluating the performance of the MPANF using real world data. We first present the experimental setup, including the characteristics of the datasets, the baselines, and the evaluation metrics. We subsequently present and analyze the experimental results, which cover a comparative assessment of out-of-sample performance, and a validation of the retrospective condition.
3.1 Experimental Setup
3.1.1 Datasets and Preprocessing
Because the MPANF’s effectiveness depends on the quality of the given movement prediction, data selection should account for both the target series and the corresponding movement prediction availability. Given the well-documented strong co-movement between the Financial Times Stock Exchange (FTSE) index and U.S. stock prices (Sarwar,, 2020; Gupta et al.,, 2024), together with the fact that the FTSE open price is recorded approximately 13 hours before the U.S. market close price on the same date, we selected eight U.S. financial time series as the target series, including Apple (AAPL), Boeing (BA), Google (GOOGL), Goldman Sachs (GS), Halliburton (HAL), Brent crude oil (OIL), Schlumberger (SLB), and Tesla (TSLA), on the basis of their strong predictive alignment with FTSE movements. Accordingly, the FTSE open price is used as the exogenous variable for generating movement predictions. The details of the movement prediction method using an exogenous variable are provided in Appendix C.
Historical data for the FTSE open price and the close prices of the eight target series were obtained from Yahoo Finance, covering the period from October 8, 2014, to September 13, 2024. The FTSE data were aligned with each target series on a day-by-day basis, with extra observations removed and missing values imputed via forward filling (Junninen et al.,, 2004). After data alignment, all time series were standardized to a uniform length of 2,500 observations and then split into in-sample and out-of-sample sets at a 1:1 ratio, ensuring sufficient data for both model estimation and evaluation. The close prices of the eight target variables and the FTSE open price are illustrated in Figures 2 and 3, respectively, with a brief description provided in Table 1. All eight target series achieved in-sample movement prediction accuracies above 0.55, suggesting that the directional signals from the FTSE opening price carried useful predictive information. For each target series, the out-of-sample forecasts by the MPANF were produced via Algorithm 1.
Figure 2: Time series of eight target variables.
Figure 3: Open price of FTSE.
Table 1: Statistics of nine financial time series.
Time Series Count Min Median Max $\mathbf{ACC}_{\text{in}}$ $\mathbf{\bar{\epsilon}}_{\text{in}}$
AAPL (Close price) 2500 20.72 54.16 234.55 0.5536 0.3842
BA (Close price) 2500 95.01 195.92 430.30 0.5568 2.5867
GOOGL (Close price) 2500 24.79 61.44 190.93 0.5504 0.4624
GS (Close price) 2500 117.58 202.27 510.25 0.5504 1.9030
HAL (Close price) 2500 4.32 33.30 151.09 0.5680 0.5252
OIL (Close price) 2500 19.33 65.20 127.98 0.5576 0.9067
SLB (Close price) 2500 11.02 49.49 74.91 0.5536 0.6596
TSLA (Close price) 2500 9.58 23.72 409.97 0.5536 0.3534
FTSE (Open price) 2500 4993.90 7210.00 8445.80
3.1.2 Baselines and Evaluation Metrics
As improvements over naive forecasts in financial time series forecasting can be difficult to achieve, even with sophisticated approaches such as machine learning and deep learning, we adopt simple and widely recognized baselines, including the naive forecast, naive forecast with drift, IMA(1,1), and linear regression, for performance comparison Rogers, (1995); Jadevicius and Huston, (2015); Franses, (2020). In particular, the linear regression model, which acts as a meta-learner to combine the naive forecast and the movement prediction, uses only the current observed value and the predicted movement as predictors.
The performance of each method was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without overpenalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows:
$RMSE=\sqrt{\frac{1}{N}\sum\limits_{i=1}^{N}{{{({{y}_{i}}-{{{\hat{y}}}_{i}})}^{2}}}},$ (16)
$MAE=\frac{1}{N}\sum\limits_{i=1}^{N}{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|},$ (17)
$MAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\left|\frac{{{y}_{i}}-{{{\hat{y}}}_{i}}}{{{y}_{i}}}\right|},$ (18)
$sMAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\frac{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|}{\frac{\left|{{y}_{i}}\right|+\left|{{{\hat{y}}}_{i}}\right|}{2}}},$ (19)
where $N$ denotes the size of the out-of-sample set, ${\hat{y}}_{i}$} denotes the predicted value, and ${y}_{i}$} denotes the actual value.
3.2 Experimental Results
3.2.1 Out-of-Sample Performance
The out-of-sample forecasting results for the eight financial time series are illustrated in Figure 4. Given that each out-of-sample set consists of 1250 time steps and that predictions from different methods visually overlap to a large extent, only the first five predicted points from each set are displayed for clarity. Small differences between the MPANF and the other forecasting methods are observed because the overall in-sample movement prediction accuracy for each time series is just slightly above 0.55, leading to a small coefficient $\alpha$, which means that the magnitude of adjustment for each time step is approximately one tenth of the base magnitude. The naive forecast, naive forecast with drift, and IMA(1,1) forecasts yield similar predictions, reinforcing the suitability of these methods as baselines and the robustness of the naive forecast as the benchmark for time series forecasting. Moreover, because the linear regression model closely mirrors the structure of the MPANF, its performance is similarly dependent on movement prediction accuracy and therefore remains close to that of the MPANF.
Figure 4: Out-of-sample forecasts by different methods.
Tables 2, 3, 4, and 5 present performance comparisons between the MPANF and the baseline methods for eight out-of-sample sets, reported in terms of the RMSE, MAE, MAPE, and sMAPE, respectively. Notably, the MPANF consistently outperforms the naive forecast, with the only exceptions being the MAPE and sMAPE on the GOOGL dataset. Specifically, the MPANF achieves the lowest RMSE in 4 out of 8 cases, the lowest MAE in 6 out of 8 cases, the lowest MAPE in 6 out of 8 cases, and the lowest sMAPE in 5 out of 8 cases, demonstrating its overall superiority over the baseline methods.
Table 2: RMSE comparison across forecasting methods for selected stocks.
Stock Naive Naive with Drift IMA(1,1) LR MPANF
AAPL 2.5937 2.5925 2.5979 2.5942 2.5931
BA 5.9080 5.9187 5.9166 5.8932 5.8829
GOOGL 2.2093 2.2085 2.2108 2.2114 2.2088
GS 5.3671 5.3660 5.3727 5.4039 5.3485
HAL 0.7334 0.7341 0.7332 0.7261 0.7271
OIL 1.8909 1.8912 1.8976 1.8940 1.8864
SLB 0.9401 0.9410 0.9401 0.9331 0.9343
TSLA 8.1524 8.1525 8.1677 8.3351 8.1515
Table 3: MAE comparison across forecasting methods for selected stocks.
Stock Naive Naive with Drift IMA(1,1) LR MPANF
AAPL 1.8950 1.8935 1.8975 1.8958 1.8943
BA 3.9691 3.9821 3.9750 3.9657 3.9533
GOOGL 1.5680 1.5660 1.5727 1.5746 1.5674
GS 3.9743 3.9737 3.9811 4.0021 3.9534
HAL 0.5423 0.5429 0.5439 0.5376 0.5368
OIL 1.2959 1.2981 1.3000 1.3073 1.2903
SLB 0.6901 0.6910 0.6907 0.6852 0.6843
TSLA 5.5732 5.5733 5.5775 5.8155 5.5722
Table 4: MAPE (%) comparison across forecasting methods for selected stocks.
Stock Naive Naive with Drift IMA(1,1) LR MPANF
AAPL 1.4081 1.4069 1.4065 1.4079 1.4074
BA 2.0601 2.0675 2.0633 2.0585 2.0514
GOOGL 1.4357 1.4340 1.4380 1.4406 1.4359
GS 1.4037 1.4036 1.4046 1.4074 1.3967
HAL 2.3517 2.3532 2.3667 2.3277 2.3216
OIL 1.8379 1.8409 1.8430 1.8456 1.8296
SLB 2.1522 2.1532 2.1574 2.1355 2.1291
TSLA 2.9511 2.9513 2.9549 3.0534 2.9509
Table 5: sMAPE (%) comparison across forecasting methods for selected stocks.
Stock Naive Naive with Drift IMA(1,1) LR MPANF
AAPL 1.4088 1.4072 1.4072 1.4088 1.4081
BA 2.0542 2.0591 2.0591 2.0500 2.0457
GOOGL 1.4352 1.4331 1.4375 1.4410 1.4354
GS 1.4035 1.4032 1.4045 1.4090 1.3966
HAL 2.3377 2.3427 2.3537 2.3051 2.3084
OIL 1.8290 1.8328 1.8336 1.8377 1.8210
SLB 2.1465 2.1503 2.1521 2.1260 2.1240
TSLA 2.9518 2.9520 2.9554 3.0767 2.9515
In some cases, baseline methods such as naive forecasting with drift, IMA(1,1), or linear regression have demonstrated superior performance, yielding lower error metric values than other models do. For example, these methods produced lower RMSE values for AAPL, GOOGL, HAL, and OIL, and lower MAEs, MAPEs, and sMAPEs for AAPL and GOOGL. However, the performances of these methods are inconsistent; they frequently fall short of the naive forecast in other cases.
Compared with the naive forecast, linear regression may sometimes perform better, particularly when movement prediction is used as an input feature. However, this occasional improvement lacks a clear theoretical justification. Its performance generally falls short of the MPANF because of a higher susceptibility to overfitting. While linear regression can fit in-sample data closely, its out-of-sample performance degrades significantly if the accuracy of the movement prediction decreases. In contrast, the MPANF’s coefficient determination is based on theoretically derived optimal conditions, rather than numerical minimization of in-sample errors. This theoretical guidance acts as a built-in regularizer, which effectively mitigates overfitting and improves the method’s ability to generalize to new and unseen data.
As shown in Figures 5, 6, 7, and 8, the MPANF outperforms the baseline methods in both accuracy and robustness across the four evaluation metrics. These empirical results align with our theoretical analysis and confirm the MPANF’s consistent improvement over the naive forecast. This superior performance, even in challenging financial scenarios characterized by heteroscedasticity and high volatility, highlights the method’s broad and reliable applicability.
Figure 5: Comparison of the RMSEs of different methods.
Figure 6: Comparison of the MAEs of different methods.
Figure 7: Comparison of the MAPEs of different methods.
Figure 8: Comparison of the sMAPEs of different methods.
3.2.2 Retrospective Validation
We further examined whether the condition expressed by Eq. (15), which specifies when the MPANF is expected to outperform the naive baseline in terms of MSE or RMSE, was satisfied across the eight target series. The purpose of this validation is to determine whether the theoretical conditions for superior performance are consistent with the actual results observed in the out-of-sample forecasts.
Table 6 summarizes the validation results for each target series. Across the eight series that exhibited improved out-of-sample RMSEs, the retrospective condition was met in all the cases. Moreover, the coefficient between the in-sample and out-of-sample periods decreased in six out of eight cases, confirming the robustness of the MPANF even under the decline in movement prediction accuracies for out-of-sample forecasting.
Table 6: Retrospective validation for eight target series.
Time Series $\hat{\alpha}^{*}_{\text{out}}$ $\bar{\epsilon}_{\text{out}}$ $\hat{\alpha}^{*}_{\text{in}}$ $\bar{\epsilon}_{\text{in}}$ Eq. (15) Satisfied?
AAPL 0.0288 1.8963 0.1072 0.3842 6
BA 0.0880 3.9675 0.1136 2.5867
GOOGL 0.0240 1.5670 0.1008 0.4624
GS 0.1248 3.9736 0.1008 1.9030
HAL 0.1248 0.5426 0.1360 0.5252
OIL 0.0880 1.2960 0.1152 0.9067
SLB 0.1232 0.6905 0.1072 0.6596
TSLA 0.0288 5.5773 0.1072 0.3534
4 Discussion
The implications of the findings and the potential limitations of the study are summarized in four points as follows.
First, in many real-world forecasting applications, decision-making requires the integration of different types of forecasts, such as binary labels and continuous values. Such integration can be technically challenging. The MPANF offers a practical solution by converting binary movement predictions into continuous point forecasts, thereby facilitating more informed decisions in diverse forecasting scenarios. The formulation of the MPANF also implies the value of incorporating directional information into a naive forecast and suggests that, in addition to developing complex regression models, we can instead develop a movement prediction-based model for accurate point forecasting.
Second, the coefficient or scaling factor in the MPANF is derived analytically on the basis of the assumption that the movement prediction is unbiased. However, although the MPANF imposes no restrictions on the choice of movement prediction models, this assumption can be violated in practice, especially when the movement prediction models that are prone to overfitting suffer substantial degradation in out-of-sample performance. Such a violation could lead to a substantial gap between the analytical estimate and the true optimum for the coefficient, reducing the practical advantages of the MPANF over the naive baseline. Moreover, generating consistent and reliable movement predictions across different segments of a financial time series remains a challenging task. Consequently, we cannot define a threshold of movement prediction accuracy for the effectiveness of MPANF, and its applicability is conditional on the availability of a meaningful and robust movement prediction for the target variable.
Third, the retrospective condition formalized in Proposition 3 offers more than a binary assessment of outperformance; it provides meaningful interpretability and practical insights into the method’s behavior. Specifically, it reveals that the MPANF’s predictive gains over the naive baseline occur only when the product of its out-of-sample movement prediction accuracy and the mean absolute increment is sufficiently large, highlighting the crucial interplay between movement prediction and increment magnitude for point forecasting performance. In addition, it serves as a valuable diagnostic tool: if the MPANF fails to improve upon the naive baseline, one can examine whether the Eq. (15) was violated, thereby distinguishing between method inadequacy and inherent unpredictability in the out-of-sample period. Furthermore, it encourages practices that enhance the stability of movement prediction by suggesting that optimal performance is approached when in-sample estimates align closely with their out-of-sample counterparts. Thus, while the condition is retrospective in nature, its implications are forward-looking, informing both interpretation and future method tuning.
Finally, while the empirical results show that even relatively modest directional signals can enhance point forecasts, the magnitude of such improvements is often marginal and context dependent. These improvements do not necessarily translate into economic benefits, especially when real-world factors such as transaction costs, slippage, and market frictions are taken into account. Trading decisions can be made after the MPANF’s performance under higher movement prediction accuracies is examined.
5 Conclusions
This study introduces the MPANF, a forecast combination method that adjusts the naive forecast by incorporating the movement prediction of the target variable. The proposed method demonstrates that systematic improvement over the naive baseline becomes theoretically and empirically possible when movement predictions are accurate and reliable. Under these conditions, movement prediction introduces genuine directional information that the naive forecast alone cannot provide. Thus, MPANF reframes the long-standing belief in the naive baseline’s supremacy: its apparent "unbeatability" in financial time series forecasting is not due to inherent optimality, but rather a consequence of the scarcity of reliable directional information. The naive forecast’s dominance, therefore, is a conditional property, not a universal one.
Several avenues for future research emerge from this work. First, because the present evaluation was conducted under modest directional accuracy, developing more advanced movement prediction models and assessing MPANF performance at higher accuracy levels could enhance its practical value in real-world trading. Second, applying MPANF to other time series domains, particularly those where the naive forecast is hard to beat yet movement prediction is feasible, could broaden its applicability. Third, extending the MPANF beyond the univariate case would address the many forecasting tasks that arise in high-dimensional settings. Finally, while this study employs a fixed adjustment for simplicity, exploring adaptive adjustment schemes may further improve the forecasting accuracy.
Declarations
  • Funding
    This study did not receive any specific grants from funding agencies in the public, commercial, or not-for-profit sectors.
  • Conflict of interest/Competing interests
    The author reports that there are no competing interests to declare.
  • Data availability
    The data are available at the following URL:https://github.com/Zhang-Cheng-76200/MPANF
  • Code availability
    The code has been made available at the following URL:https://github.com/Zhang-Cheng-76200/MPANF
  • Author contributions
    The author confirms being the sole contributor of this work.
Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in}}$
According to Eq. (2), the in-sample MSE of the naive forecast, denoted by $MSE_{\text{in}}^{\text{Naive}}$, can be expressed as:
$MSE_{\text{in}}^{\text{Naive}}=\frac{1}{N}\left(\sum_{t\in\text{in}}\left(y_{t}-y_{t-1}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{in}}\left((y_{t-1}+d_{t}\cdot|\epsilon_{t}|)-y_{t-1}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$ (A1)
By using Eq. (2) and Eq. (4) to replace $y_{t}$ and $y_{t-1}$ in Eq. (A1), respectively, the in-sample MSE of the MPANF, denoted by $MSE_{\text{in}}^{\text{MPANF}}$, can be expressed as follows:
$MSE_{\text{in}}^{\text{MPANF}}=\frac{1}{N}\left(\sum_{t\in\text{in}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+d_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{d}_{t}\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}})\right]^{2}\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left[(y_{t-1}+d_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{d}_{t}\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}})\right]^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (A2)
The in-sample MSE difference between the naive forecast and the MPANF, denoted by $\Delta MSE_{\text{in}}$, can subsequently be derived via Eqs. (A1) and (A2):
$\Delta MSE_{\text{in}}=MSE_{\text{in}}^{\text{Naive}}-MSE_{\text{in}}^{\text{MPANF}}$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}-\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)-\sum_{t\in\text{incorrect}}\left(2\cdot\alpha_{\text{in}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (A3)
Assuming that the in-sample movement prediction is unbiased and contains sufficient numbers of both correct and incorrect predictions, the following approximations hold on the basis of the law of large numbers (LLN):
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in}}$ (A4)
Therefore, Eq. (A3) can be simplified as follows::
$\Delta MSE_{\text{in}}=MSE_{\text{in}}^{\text{Naive}}-MSE_{\text{in}}^{\text{MPANF}}$
$\approx\frac{1}{N}\left(2\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}}\cdot(n_{\text{in}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in}})-n_{\text{in}}^{\text{correct}}\cdot\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}-2\cdot\alpha_{\text{in}}\cdot\bar{\epsilon}_{\text{in}}\cdot(n_{\text{in}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in}})-n_{\text{in}}^{\text{incorrect}}\cdot\alpha_{\text{in}}^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)$
$=\left(4\cdot\alpha_{\text{in}}\cdot ACC_{\text{in}}-\alpha_{\text{in}}^{2}-2\cdot\alpha_{\text{in}}\right)\cdot\bar{\epsilon}_{\text{in}}^{2}$ (A5)
Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out}}$
When $\hat{\alpha}_{\text{in}}^{*}$ and $\bar{\epsilon}_{\text{in}}$ are adopted by the MPANF on the out-of-sample set, the out-of-sample MSE difference between the naive forecast and the MPANF, denoted by $\Delta MSE_{\text{out}}$, can be expressed as follows:
$\Delta MSE_{\text{out}}=MSE_{\text{out}}^{\text{Naive}}-MSE_{\text{out}}^{\text{MPANF}}$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}-\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)-\sum_{t\in\text{incorrect}}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (B1)
Similarly, assuming that the out-of-sample movement prediction is unbiased and contains sufficient numbers of both correct and incorrect predictions, the following approximations hold on the basis of the LLN:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out}}$ (B2)
Therefore, Eq. (B1) can further be expressed as:
$\Delta MSE_{\text{out}}=MSE_{\text{out}}^{\text{Naive}}-MSE_{\text{out}}^{\text{MPANF}}$
$\approx\frac{1}{M}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot(m_{\text{out}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out}})\right.-m_{\text{out}}^{\text{correct}}\cdot(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\left.-2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot(m_{\text{out}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out}})-m_{\text{out}}^{\text{incorrect}}\cdot(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)$
$=2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot ACC_{\text{out}}\cdot\bar{\epsilon}_{\text{out}}-2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot(1-ACC_{\text{out}})\cdot\bar{\epsilon}_{\text{out}}-(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}$
$\approx 2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot\hat{\alpha}_{\text{out}}^{*}\cdot\bar{\epsilon}_{\text{out}}-(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}$ (B3)
Appendix C. Movement Prediction with an Exogenous Variable
A simple but effective method for generating movement prediction involves leveraging the co-movement between the target variable and an exogenous variable. The selection of the exogenous variable is typically guided by domain knowledge, and the movement prediction of the target variable, denoted by $\hat{d}_{t}^{\text{target}}$, can be simply obtained as:
$\hat{d}_{t}^{\text{target}}=\hat{d}_{t}^{\text{exo}},$ (C1)
where $\hat{d}_{t}^{\text{exo}}$ represents the movement prediction of the exogenous variable at time step $t$. In particular, if the value of the exogenous variable is recorded sufficiently ahead of the target variable for the same time step, such as in financial markets operating in different time zones, the actual movement of the exogenous variable can directly serve as the movement prediction for the target variable:
$\hat{d}_{t}^{\text{target}}=d_{t}^{\text{exo}}=\begin{cases}1,&\text{if }y_{t}^{\text{exo}}>y_{t-1}^{\text{exo}};\\ -1,&\text{if }y_{t}^{\text{exo}}\leq y_{t-1}^{\text{exo}};\end{cases},$ (C2)
where $d_{t}^{\text{exo}}$ is the actual movement of the target variable at time step $t$. $y_{t}^{\text{exo}}$ and $y_{t-1}^{\text{exo}}$ denotes the actual values of the exogenous variable at time steps $t$ and t1t-1, respectively.